# Detecting Measurement-Induced Entanglement Transitions With Unitary Mirror Circuits

### v1

@@ II.1 Extremizing the logand
$\begin{split}F^{\pm}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma,p_{\min},p_{\max}}}\right)&=\pm\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\max_{\begin{subarray}{c}\{p_{1},\dotsc,p_{m}\}\\ \sum_{i=1}^{m}\mu_{i}=\Sigma\\ p_{\min}\leq\mu_{i}\leq p_{\max}\end{subarray}}\mathopen{}\mathclose{{}\left[{\pm\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m}\mu_{i}^{\alpha}}}\right],\end{split}$ (S7)
@@
$F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma,0,p_{\max}}}\right)=\lfloor\frac{\Sigma}{p_{\max}}\rfloor(p_{\max})^{\alpha}+\mathopen{}\mathclose{{}\left({\Sigma-\lfloor\frac{\Sigma}{p_{\max}}\rfloor p}}\right)^{\alpha},$ (S8b)
@@ II.1.1 Proof S8a
$\begin{split}F^{+}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,p_{\min},p_{\max}}}\right)&=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\max_{\begin{subarray}{c}\{p,\dotsc,p_{m+1}\}\\ \sum_{i=1}^{m+1}\mu_{i}=\Sigma\\ p_{\min}\leq\mu_{i}\leq p_{\max}\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m+1}\mu_{i}^{\alpha}}}\right]\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\max_{\begin{subarray}{c}p\\ \frac{\Sigma}{m+1}\leq p\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]p^{\alpha}+\max_{\begin{subarray}{c}\{p,\dotsc,p_{m}\}\\ \sum_{i=1}^{m}\mu_{i}=\Sigma-p\\ p_{\min}\leq\mu_{i}\leq p\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m}\mu_{i}^{\alpha}}}\right]}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\max_{\begin{subarray}{c}p\\ \frac{\Sigma}{m+1}\leq p\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({p^{\alpha}+F^{+}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma-p,p_{\min},p}}\right)}}\right)}\Big{]}.\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\max_{\begin{subarray}{c}p\\ \frac{\Sigma}{m+1}\leq p\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({p^{\alpha}+m\mathopen{}\mathclose{{}\left[{\frac{\Sigma-p}{m}}}\right]^{\alpha}}}\right)}\Big{]}.\end{split}$
@@
$\begin{gathered}\frac{\partial}{\partial p}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({p^{\alpha}+m\mathopen{}\mathclose{{}\left[{\frac{\Sigma-p}{m}}}\right]^{\alpha}}}\right)}}\right]=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\alpha\mathopen{}\mathclose{{}\left({p^{\alpha-1}-\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({\Sigma-p}}\right)/m}}\right]^{\alpha-1}}}\right)<0\quad\text{for }p>\Sigma/(m+1),\\ \Rightarrow F^{+}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,p_{\min},p_{\max}}}\right)=\mathopen{}\mathclose{{}\left({m+1}}\right)\mathopen{}\mathclose{{}\left({\frac{\Sigma}{m+1}}}\right)^{\alpha}.\end{gathered}$
@@ II.1.2 Proof S8b
Assume $F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma,0,p_{\max}}}\right)=\lfloor\frac{\Sigma}{p_{\max}}\rfloor(p_{\max})^{\alpha}+\mathopen{}\mathclose{{}\left({\Sigma-\lfloor\frac{\Sigma}{p_{\max}}\rfloor p}}\right)^{\alpha}$. Then
@@
$\begin{split}F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,0,p_{\max}}}\right)&=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\min_{\begin{subarray}{c}\{p,\dotsc,p_{m+1}\}\\ \sum_{i=1}^{m+1}\mu_{i}=\Sigma\\ 0\leq\mu_{i}\leq p_{\max}\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m+1}\mu_{i}^{\alpha}}}\right]\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}p\\ \frac{\Sigma}{m+1}\leq p\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]p^{\alpha}+\min_{\begin{subarray}{c}\{p,\dotsc,p_{m}\}\\ \sum_{i=1}^{m}\mu_{i}=\Sigma-p\\ 0\leq\mu_{i}\leq p\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m}\mu_{i}^{\alpha}}}\right]}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}p\\ \frac{\Sigma}{m+1}\leq p\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({p^{\alpha}+F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma-p,0,p}}\right)}}\right)}\Big{]}.\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}p\\ \frac{\Sigma}{m+1}\leq p\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({p^{\alpha}+\lfloor\frac{\Sigma-p}{p}\rfloor p^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-p-\lfloor\frac{\Sigma-p}{p}\rfloor p}}\right]^{\alpha}}}\right)}\Big{]}.\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}p\\ \frac{\Sigma}{m+1}\leq p\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\lfloor\frac{\Sigma}{p}\rfloor p^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\lfloor\frac{\Sigma}{p}\rfloor p}}\right]^{\alpha}}}\right)}\Big{]}.\end{split}$
Minimizing,
@@
$\begin{split}\frac{\partial}{\partial p}&\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\lfloor\frac{\Sigma}{p}\rfloor p^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\lfloor\frac{\Sigma}{p}\rfloor p}}\right]^{\alpha}}}\right)}}\right]_{\Sigma/p\notin\mathbb{N}}=\\ &\quad\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left[{\alpha\lfloor\frac{\Sigma}{p}\rfloor p^{\alpha-1}\mathopen{}\mathclose{{}\left({1-\mathopen{}\mathclose{{}\left[{\frac{\Sigma}{p}-\lfloor\frac{\Sigma}{p}\rfloor}}\right]^{\alpha-1}}}\right)}}\right]<0\\ &\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\lfloor\frac{\Sigma}{p}\rfloor p^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\lfloor\frac{\Sigma}{p}\rfloor p}}\right]^{\alpha}}}\right)}}\right]_{p=\Sigma/\ell+\varepsilon}-\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\lfloor\frac{\Sigma}{p}\rfloor p^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\lfloor\frac{\Sigma}{p}\rfloor p}}\right]^{\alpha}}}\right)}}\right]_{p=\Sigma/\ell}=\\ &\quad\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({\ell-1}}\right)\mathopen{}\mathclose{{}\left({\Sigma/\ell+\varepsilon}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\mathopen{}\mathclose{{}\left({\ell-1}}\right)\mathopen{}\mathclose{{}\left({\frac{\Sigma}{\ell}+\varepsilon}}\right)}}\right]^{\alpha}-\ell\mathopen{}\mathclose{{}\left({\Sigma/\ell}}\right)^{\alpha}}}\right]\\ &\quad=-\alpha\mathopen{}\mathclose{{}\left\lvert 1-\alpha}\right\rvert\mathopen{}\mathclose{{}\left({\Sigma/\ell}}\right)^{\alpha-2}\ell\mathopen{}\mathclose{{}\left({\ell-1}}\right)\varepsilon^{2}/2+O\mathopen{}\mathclose{{}\left({\varepsilon}}\right)^{3}<0\end{split}$
@@
$\begin{split}F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,0,p_{\max}}}\right)&=\lfloor\frac{\Sigma}{p_{\max}}\rfloor(p_{\max})^{\alpha}+\mathopen{}\mathclose{{}\left({\Sigma-\lfloor\frac{\Sigma}{p_{\max}}\rfloor p_{\max}}}\right)^{\alpha}\end{split}$

### v2

@@ II.1 Extremizing the logand
$\begin{split}F^{\pm}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma,p_{\min},p_{\max}}}\right)&=\pm\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\max_{\begin{subarray}{c}\{\mu_{1},\dotsc,\mu_{m}\}\\ \text{so that}\\ \sum_{i=1}^{m}\mu_{i}=\Sigma\\ \forall\mu_{i},\;p_{\min}\leq\mu_{i}\leq p_{\max}\end{subarray}}\mathopen{}\mathclose{{}\left[{\pm\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m}\mu_{i}^{\alpha}}}\right],\end{split}$ (S7)
@@
$F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma,0,p_{\max}}}\right)=\lfloor\frac{\Sigma}{p_{\max}}\rfloor(p_{\max})^{\alpha}+\mathopen{}\mathclose{{}\left({\Sigma-\lfloor\frac{\Sigma}{p_{\max}}\rfloor p_{\max}}}\right)^{\alpha},$ (S8b)
@@ II.1.1 Proof S8a
$\begin{split}F^{+}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,p_{\min},p_{\max}}}\right)&=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\max_{\begin{subarray}{c}\{\mu_{1},\dotsc,\mu_{m+1}\}\\ \sum_{i=1}^{m+1}\mu_{i}=\Sigma\\ p_{\min}\leq\mu_{i}\leq p_{\max}\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m+1}\mu_{i}^{\alpha}}}\right]\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\max_{\begin{subarray}{c}{\bar{\mu}}\\ \frac{\Sigma}{m+1}\leq{\bar{\mu}}\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]{\bar{\mu}}^{\alpha}+\max_{\begin{subarray}{c}\{\mu_{2},\dotsc,\mu_{m+1}\}\\ \sum_{i=2}^{m+1}\mu_{i}=\Sigma-{\bar{\mu}}\\ p_{\min}\leq\mu_{i}\leq{\bar{\mu}}\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=2}^{m+1}\mu_{i}^{\alpha}}}\right]}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\max_{\begin{subarray}{c}{\bar{\mu}}\\ \frac{\Sigma}{m+1}\leq{\bar{\mu}}\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha}+F^{+}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma-{\bar{\mu}},p_{\min},{\bar{\mu}}}}\right)}}\right)}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\max_{\begin{subarray}{c}{\bar{\mu}}\\ \frac{\Sigma}{m+1}\leq{\bar{\mu}}\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha}+m\mathopen{}\mathclose{{}\left[{\frac{\Sigma-{\bar{\mu}}}{m}}}\right]^{\alpha}}}\right)}\Big{]}.\end{split}$
Here, we separated the term in the brackets into the largest coefficient $\bar{\mu}$ and the rest, and then identified the remaining sum as the maximized function for $m$, and inserted its value as per the induction assumption.
@@
$\begin{gathered}\frac{\partial}{\partial{\bar{\mu}}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha}+m\mathopen{}\mathclose{{}\left[{\frac{\Sigma-{\bar{\mu}}}{m}}}\right]^{\alpha}}}\right)}}\right]=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\alpha\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha-1}-\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({\Sigma-{\bar{\mu}}}}\right)/m}}\right]^{\alpha-1}}}\right)<0\quad\text{for }{\bar{\mu}}>\Sigma/(m+1),\\ \Rightarrow F^{+}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,p_{\min},p_{\max}}}\right)=\mathopen{}\mathclose{{}\left({m+1}}\right)\mathopen{}\mathclose{{}\left({\frac{\Sigma}{m+1}}}\right)^{\alpha}.\end{gathered}$
@@ II.1.2 Proof S8b
Assume $F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma,0,p_{\max}}}\right)=\lfloor\frac{\Sigma}{p_{\max}}\rfloor(p_{\max})^{\alpha}+\mathopen{}\mathclose{{}\left({\Sigma-\lfloor\frac{\Sigma}{p_{\max}}\rfloor p}}\right)^{\alpha}$. Then, repeating the same procedure,
@@
$\begin{split}F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,0,p_{\max}}}\right)&=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\min_{\begin{subarray}{c}\{\mu_{1},\dotsc,\mu_{m+1}\}\\ \sum_{i=1}^{m+1}\mu_{i}=\Sigma\\ 0\leq\mu_{i}\leq p_{\max}\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m+1}\mu_{i}^{\alpha}}}\right]\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}{\bar{\mu}}\\ \frac{\Sigma}{m+1}\leq{\bar{\mu}}\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]{\bar{\mu}}^{\alpha}+\min_{\begin{subarray}{c}\{\mu_{2},\dotsc,\mu_{m+1}\}\\ \sum_{i=2}^{m+1}\mu_{i}=\Sigma-{\bar{\mu}}\\ 0\leq\mu_{i}\leq{\bar{\mu}}\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=2}^{m+1}\mu_{i}^{\alpha}}}\right]}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}{\bar{\mu}}\\ \frac{\Sigma}{m+1}\leq{\bar{\mu}}\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha}+F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma-{\bar{\mu}},0,{\bar{\mu}}}}\right)}}\right)}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}{\bar{\mu}}\\ \frac{\Sigma}{m+1}\leq{\bar{\mu}}\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha}+\lfloor\frac{\Sigma-{\bar{\mu}}}{{\bar{\mu}}}\rfloor{\bar{\mu}}^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-{\bar{\mu}}-\lfloor\frac{\Sigma-{\bar{\mu}}}{{\bar{\mu}}}\rfloor{\bar{\mu}}}}\right]^{\alpha}}}\right)}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}{\bar{\mu}}\\ \frac{\Sigma}{m+1}\leq{\bar{\mu}}\leq p_{\max}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}}}\right]^{\alpha}}}\right)}\Big{]}.\end{split}$
Minimizing this, for any integer $\ell$,
@@
$\begin{split}\frac{\partial}{\partial{\bar{\mu}}}&\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}}}\right]^{\alpha}}}\right)}}\right]_{\Sigma/{\bar{\mu}}\notin\mathbb{N}}=\\ &\quad\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left[{\alpha\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}^{\alpha-1}\mathopen{}\mathclose{{}\left({1-\mathopen{}\mathclose{{}\left[{\frac{\Sigma}{{\bar{\mu}}}-\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor}}\right]^{\alpha-1}}}\right)}}\right]<0\\ &\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}}}\right]^{\alpha}}}\right)}}\right]_{{\bar{\mu}}=\Sigma/\ell+\varepsilon}-\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\lfloor\frac{\Sigma}{{\bar{\mu}}}\rfloor{\bar{\mu}}}}\right]^{\alpha}}}\right)}}\right]_{{\bar{\mu}}=\Sigma/\ell}=\\ &\quad\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({\ell-1}}\right)\mathopen{}\mathclose{{}\left({\Sigma/\ell+\varepsilon}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left[{\Sigma-\mathopen{}\mathclose{{}\left({\ell-1}}\right)\mathopen{}\mathclose{{}\left({\frac{\Sigma}{\ell}+\varepsilon}}\right)}}\right]^{\alpha}-\ell\mathopen{}\mathclose{{}\left({\Sigma/\ell}}\right)^{\alpha}}}\right]\\ &\quad=-\alpha\mathopen{}\mathclose{{}\left\lvert 1-\alpha}\right\rvert\mathopen{}\mathclose{{}\left({\Sigma/\ell}}\right)^{\alpha-2}\ell\mathopen{}\mathclose{{}\left({\ell-1}}\right)\varepsilon^{2}/2+O\mathopen{}\mathclose{{}\left({\varepsilon}}\right)^{3}<0\end{split}$
@@
$\begin{split}F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,0,p_{\max}}}\right)&=\lfloor\frac{\Sigma}{p_{\max}}\rfloor(p_{\max})^{\alpha}+\mathopen{}\mathclose{{}\left({\Sigma-\lfloor\frac{\Sigma}{p_{\max}}\rfloor p_{\max}}}\right)^{\alpha}.\end{split}$